Prev Next

SW&IT Meeting Minutes 2021-08-09

Attendees: David, Matthieu, Michael, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210726
LR20210726 - Bec, Matthieu 40.0
LR20210726 - Edmeades, David 0.0
LR20210726 - Grenz, Paul 37.0
LR20210726 - Hooper, Stephen 0.0
LR20210726 - Zhang, Yang 80.0
LR20210726 - Zhang, Xianyu 40.0

News from management

  • Brandon + Glenn starting next week Aug 16.

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • MODS KVM upgrade
  • obs 5/6 readiness
  • support schedule 2021B

Team reports

MB

  • Last week
    • OC, Management, ATT PM
    • cobbler ks centos6 template, dms.tucson f32
    • on-boarding preps Glenn, Brandon
    • AO SX s/w
  • This week
    • on-boarding preps Glenn, Brandon
    • dms.tucson iterations
    • on-boarding preps Glenn, Brandon
    • AO SX s/w

YZ

  • Previous:
    • Worked with Paul on TCS and Argos light source debugging and testing.
    • Updated the TCS releasing and patching document.
    • Some MODS IDL issues (should have resolved) with Olga.
    • Upgrading OSCO background computational process.
    • Reviewing some TMS documents.
  • Next:
    • Continue to assist Paul with any testing and debugging plans.
    • Finish upgrading OSCO background process.
    • MODS new KVM installation (and testing perhaps).
    • Initiate some MODS related discussions with OSU (e.g., KVM, software management, MODS IDL, WCS header info plans).
    • IT requests (Fedora VM /lbt mount).
    • Investigate some MODS WCS header related information.

DE

  • Followup on mountain Synology disk replacement; should be on its way
  • Autodesk licensing; Autodesk is having issues with their system so I haven't been able to generate a license file, but currently not a showstopper
  • Matlab yearly license generated, no feedback yet from Mike but I don't anticapate any trouble
  • Shopping for new hires
  • Meeting today with Ketiv regarding Vault server upgrade. I'll get info about how long they anticipate it will take, transitions, etc.
  • AO telemetry stuff

PSG

  • Last week
    • Tested new interface for white light source.
    • Verified it works properly
    • On a tangent, found (and fixed) long standing bug in OSS
  • This week
    • Coordinating work with Gustavo to update ARGOS swing arm controller.

SH

  • SOUL switch in the lab
  • DELL with Louise
  • Aluminizing (Dan R., network, impromptu meetings).
  • increased disk space for robs
next:
  • LDAP on mountain
  • DNS on mountain
Topic revision: r8 - 09 Aug 2021, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback